### v1

@@ B.3 Re-computing $c/a$ but with test satellites within 300 kpc
Fig. 18 shows the evolution of $c/a$ for the $MW_{\text{fiducial}}$ potential model, comparing cases with and without the exclusion of test satellites beyond 300 kpc. The dashed green line represents the backward evolution when these satellites are excluded, while the solid green line shows the backward evolution including all test particles. The black lines indicate the forward integration. The exclusion of satellites beyond 300 kpc does not significantly alter the overall evolution of $c/a$. While removing some test satellites has a minor effect on the backward integration of $c/a$, this effect is not substantial if satellites beyond 300 kpc are excluded.

### v2

@@ B.3 Re-computing $c/a$ but with test satellites within 300 kpc
Figure 18 shows the evolution of $c/a$ for the $MW_{\text{fiducial}}$ potential model, comparing cases with and without the exclusion of test satellites beyond 300 kpc. The dashed green line represents the backward evolution when these satellites are excluded, while the solid green line shows the backward evolution including all test particles. The black lines indicate the forward integration. The exclusion of satellites beyond 300 kpc does not significantly alter the overall evolution of $c/a$. While removing some test satellites has a minor effect on the backward integration of $c/a$, this effect is not substantial if satellites beyond 300 kpc are excluded.